Steam at 15 bar and 250oC is expanded reversibly in a closed system to 5 bar. At this pressure the steam is just dry saturated. For a mass of 1 kg calculate the following.

i. The final volume.

ii. The change in internal energy.

iii. The work done.


1
Expert's answer
2021-09-07T23:58:09-0400

a)

Pv=RT

v="\frac{RT}{P}"

="\frac{0.46*523}{5}"

=48.116m3

b)

∆u=Q-W

W=mRT1ln("\frac{P_1}{P_2}" )

=1*0.46*523ln("\frac{15}{5}" )

=264.30J

Q=0

∆u=-264.30J

c)

W=264.30J